OpenStudy (anonymous):

sWrite the equation of the circle in standard form. Find the center, radius, intercepts, and graph the circle. x2+y2+8x+2y+8=0

OpenStudy (souvik):

(x^2+8x+16)+(y^2+2y+1)+8=16+1 (x+4)^2+(y+1)^2=9

OpenStudy (souvik):

so center=(-4,-1) radius=3 unit